OpenStudy (anonymous):

Find the least common denominator for the following pair of rational expressions.

OpenStudy (anonymous):

\[\left(\begin{matrix}3 \\ p^2+12p+32\end{matrix}\right),\left(\begin{matrix}2 \\ p^2+2p-8\end{matrix}\right)\]

OpenStudy (anonymous):

factorize the denominators first

OpenStudy (anonymous):

oh yeah their is supposed to be a line between top and bottom numbers

OpenStudy (anonymous):

p^2+12p+32=(p+4)(p+8) p^2+2p-8=(p-2)(p+4) so the least common denominator is:(p+8)(p+4)(p-2)

OpenStudy (anonymous):

thanks